# appendixReferences

### v3

@@ A.2.2 Proof of Theorem 4
By (44), it holds that $\hat{\Phi}_{1}^{\textup{Haj}}$ and $\hat{\Phi}_{2}^{\textup{Haj}}$ are within $O_{P}(N^{1/2})$ of $\Phi$, and hence $\min\left(\hat{\Phi}_{1}^{\textup{Haj}}\,,\,\hat{\Phi}_{2}^{\textup{Haj}}\right)$ is within $O_{P}(N^{1/2})$ as well. As $\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)$ upper bounds $\hat{\Phi}_{k}$, it follows that
$\min\left(\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\,,\,\max_{\phi\in\{0,1\}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)\geq\Phi+O_{P}(N^{1/2}),$
which implies that $\hat{\tau}\leq\tau+O_{P}(N^{1/2})$, proving (45). ∎
@@
It holds that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}^{\textup{Haj}}(\phi)$ $=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}+\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{i})$
$=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{i}+N$ (92)
=$=\begin{cases}N+\Delta^{{\textup{Haj}}}&\textup{ if }k=1\\ N-\Delta^{{\textup{Haj}}}&\textup{ if }k=2\end{cases}$ (93)
where (92) holds since the definition of $\hat{N}_{0}$ implies that
$\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}=1,$
and $\Delta^{{\textup{Haj}}}$ denotes
$\Delta^{{\textup{Haj}}}=\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}Y_{i},$
for $\hat{N}_{a}$ defined as in the statement of the theorem.
To show (46), we observe that
$\hat{\tau}$ $=N-\min\left(\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ \max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)$
$=\max\left(N-\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ N-\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right)$
$=\max\left(-\Delta^{{\textup{Haj}}},\Delta^{{\textup{Haj}}}\right)$
$=\left|\Delta^{{\textup{Haj}}}\right|$
where the inequality follows by substitution of (93). ∎
A.2.3 Proof of Theorem 5
Proof of Theorem 5, Eq. (47).
For $C_{ij}$ given by (54), let $M\in\mathbb{R}^{N\times N}$ denote the matrix given by
$M_{ij}=\begin{cases}C_{ij}\cdot\dfrac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}&\textup{ if $\phi_{i}^{*}\phi_{j}^{*}=1$}\\ 0&\textup{ else }\end{cases}$
so that $V_{k}=\sum_{ij}M_{ij}$.
By Lemma 2 and (58) of Lemma 3, $C>0$ exists such that $|M_{ij}|$ may be bounded by
$|M_{ij}|\leq\begin{cases}\frac{C\log N}{N\rho_{\min}}&i,j\in\mathcal{I}_{0}\\ \frac{C}{\rho_{\min}}&i,j\in\mathcal{I}_{1}.\end{cases}$ (94)
By Assumption 1, it holds that
$\max_{i\in[N]}|\{j:(i,j)\in\mathcal{I}_{1}\}|\leq d_{\max}.$ (95)
Combining (94) and (95) implies that each row or column of $M$ has entries summing to at most $C(\log N)/\rho_{\min}+Cd_{\max}/\rho_{\min}$.
Changing $X$ in a single entry changes at most $d_{\max}$ rows and columns of $M$. As a result, for a fixed constant $C_{2}>0$ the bounded differences condition holds for $N\geq 2$ and all $X,X^{\prime}$ differing in a single entry that
$|V_{k}(X)-V_{k}(X^{\prime})|\leq C_{2}\log N$
and it holds by McDiarmid’s inequality that
$P(|V_{k}-\mathbb{E}V_{k}|\geq\epsilon)\leq 2\exp\left(\frac{-2\epsilon^{2}}{C_{2}N(\log N)^{2}}\right)$
which implies that $|V-\mathbb{E}V_{k}|=O_{P}(N^{1/2+\delta})$ for any $\delta>0$. ∎
Proof of Theorem 5, Eq. (48).
Recall that $C_{ij}$ is given by (54). To show that $V_{k}$ is not anti-conservative, we observe that
$\mathbb{E}V_{k}$ $=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \mathbb{E}\left[\mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg{|}T_{i},T_{j}\right]\right]$
$=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg{|}T_{i},T_{j}\right]\right]$
$=\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\cdot C_{ij}(t_{i},t_{j})\cdot$
@@ A.2.3 Proof of Theorem 5
$\geq\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$ (96)
$\geq\sum_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\ {\operatorname{Cov}}(v_{j},v_{j})$ (97)
where (96) holds by Lemma 4, and (97) holds by Lemma 5. It can be seen that (97) equals the claim of (48) made by Theorem 5.
To verify the conditions for unbiasedness given by Theorem 5, we observe that if $T_{i}$ is the constant mapping for all units, then $V_{k}$ simplifies to
$V_{k}=\sum_{i:\phi^{*}=1}\sum_{j:\phi^{*}=1}{{\operatorname{Cov}}}\left(v_{i},v_{j}\right)\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)}$
As this is an importance sampled estimator of covariance terms appearing in ${\operatorname{Var}\,}\Phi_{k}^{{\textup{Haj}}}$, it is unbiased provided that the sampling probabilities $P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)$ are non-zero for each term.
A.2.4 Proof of Theorem 6
Proof of Theorem 6.
It holds that
$CI-\tau$ $=\Phi-\min\left\{\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\ ,\ \max_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$
$\leq\Phi-\min\left\{\hat{\Phi}_{1}+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \hat{\Phi}_{2}+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}$
$=\max\left\{\Phi-\hat{\Phi}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \Phi-\hat{\Phi}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}$ (98)
Let $\mathcal{E}_{k}$ denote the event
$\mathcal{E}_{k}=\left\{\Phi-\hat{\Phi}_{k}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\geq 0\right\}.$
$P(\mathcal{E}_{1}\cup\mathcal{E}_{2})\leq\alpha+o(1)$ (99)
Combining (98) and (99) yields that
$P(CI-\tau>0)\leq\alpha+o(1)$ (100)
proving that $CI\leq\tau$ with probability converging to at least $1-\alpha$. ∎
A.3 Proof of Auxilliary Lemmas
@@
For each $i\in[N]$, it can be seen that $\|u_{i}\|\leq 1/p_{\min}$ and that $u_{i}$ depends on $X$ only through $\bar{Z}_{i}$ and $T_{i}$, and hence only through $X_{\Gamma_{i}}$. As a result, changing $X$ in a single entry changes at most $d_{\max}$ entries of $u_{1},\ldots,u_{N}$, and hence changes $\bar{u}$ by at most $C=d_{\max}/p_{\min}$. Applying McDiarmid’s inequality to the $\ell$th entry of $\bar{u}$ for yields that
$P(|\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}|\geq\epsilon)\leq\exp\left(-2\epsilon^{2}/NC\right)$
and hence $\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}=O_{P}(N^{-1/2})$, and $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(N^{-1/2})$ as well. ∎
Proof of Lemma 2.
For $i,j$ satisfying $\phi_{i}^{*}\phi_{j}^{*}=1$, let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
@@
Recall $C_{ij}$ is given by (54). To show (55), we observe that for $i,j\in\mathcal{I}_{0}$ that
@@
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$={\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}},\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
@@
$\hskip 28.45274pt+\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)$ (102)
@@
To show (56), we observe for $i,j\in\mathcal{I}_{1}$ that
@@
$\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho_{\min}}$ (103)
$\leq\frac{1}{p_{\min}^{2}}\left(\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]\cdot\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{kj}})^{2}}|F_{\{i,j\}}\right]\right)^{1/2}+$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho_{\min}}$
$=O\left(1\right)+O\left(\frac{\log N}{N}\right)$ (104)
where (103) uses Lemma 6 and (57), and (104) uses Lemma 7 and also because
@@
$=O(1/N)+O(1)$
@@
Given $i,j\in[N]$, let $\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}=\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$ denote the space of possible treatment assignments for $X_{\Gamma_{i}\cup\Gamma_{j}}$. By Assumption 1, the unit treatments are independently assigned, with $p_{\min}\leq P(X_{i}=1)\leq 1-p_{\min}$. Since $|\Gamma_{i}|\leq d_{\max}$, it follows for $\rho_{\min}=(p_{\min})^{2d_{\max}}$ that
$P(X_{\Gamma_{i}\cup\Gamma_{j}}=x)\geq\rho_{\min}\qquad\textup{ for all $x\in\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$}$ (105)
As $T_{i}$ and $T_{j}$ depends on $X$ only through $X_{\Gamma_{i}}$ and $X_{\Gamma_{j}}$, we may write
$P(T_{i}=t_{i},T_{j}=t_{j})$ $=P(X_{\Gamma_{i}\cup\Gamma_{j}}\in\mathcal{S})$
for some (possibly empty) set $\mathcal{S}\subseteq\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}$. It follows that $P(T_{i}=t_{i},T_{j}=t_{j})$ equals zero if $\mathcal{S}$ is empty, and otherwise lower bounded by $\rho_{\min}$ by (105). This establishes
[…]
which equals (57).
To establish (58), we observe that similar to $T_{i}$, if $\phi_{i}^{*}=1$ then the occurrence of the event $\{\bar{Z}_{i}=Y_{ki}\}$ depends only on $X$ through $X_{\Gamma_{i}}$. It follows that
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})=P(X_{\Gamma_{i}\cup\Gamma_{j}}\in\mathcal{S})$
for some (possibly empty) set $\mathcal{S}\subseteq\mathcal{X}_{\Gamma_{i}\cup\Gamma_{j}}$. Applying (105) yields
[…]
Combining (106) with Bayes’ rule, which implies
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})\geq P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})$
with equality when $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj},T_{i},T_{j})=0$, establishes that
[…]
proving the claim. ∎
Proof of Lemma 4.
Recall that $v_{i}$ and $C_{ij}$ are respectively given by (37) and (54). If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then
@@
To show (60), we observe that if $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$ then
@@
To show (61), we observe that if $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
@@
$\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
@@
$\hat{N}_{Q}$ $=\sum_{i\in Q}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ $\hat{N}_{\bar{S}}$ $=\sum_{i\in\bar{S}}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$
@@
Under Assumption 1, it can be seen that changing $X$ in a single entry changes at $\hat{N}_{Q}$ by at most $C=d_{\max}/p_{\min}$. As a result, by McDiarmid’s inequality it holds that
$P\left(\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon\ \Big{|}\ X_{\Gamma_{S}}\right)\leq\exp\left(-\frac{\epsilon^{2}}{NC^{2}}\right)$
Thus for $\epsilon_{B}=\sqrt{3C^{2}N\cdot\log N}$ and the indicator
@@
it holds that
$P(B=1\ |\ X_{\Gamma_{S}})\leq 1/N^{3}.$ (113)
For $i,j:\phi_{j}^{*}\phi_{i}^{*}=1$, we let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that

### v4

@@ D.4 Proof of Theorem 7
For $C_{ij}$ given by (128), let $Q\in\mathbb{R}^{N\times N}$ denote the matrix given by
$Q_{ij}=\begin{cases}C_{ij}\cdot\dfrac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}&\textup{ if $i,j\in{\mathcal{I}}$}\\ 0&\textup{ else }\end{cases}$ (139)
so that $\hat{V}_{k}=\sum_{ij}Q_{ij}$.
To bound $|Q_{ij}|$, we may apply Lemma 2 and (55) to bound $|C_{ij}|$ and then (56) to bound the term $1/P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$ appearing in (139), yielding for some $C_{1}>0$ that
$|Q_{ij}|\leq\begin{cases}\displaystyle{\frac{C_{1}\log M}{M}}&i,j\in\mathcal{J}_{0}\\ \displaystyle{C_{1}}&i,j\in\mathcal{J}_{1}.\end{cases}$ (140)
By (137) it holds that
$\max_{i\in[N]}|\{j:(i,j)\in\mathcal{J}_{1}\}|\leq\frac{d^{2}CN}{M}.$ (141)
Changing $X_{G_{m}}$ changes at most $\frac{dCN}{M}$ rows and columns of $Q$, by (138). Combining (140) and (141) implies that each row or column of $Q$ has entries summing to at most $C_{2}\frac{N}{M}(\log M+1)$ for some $C_{2}>0$. It thus follows for some $C_{3}>0$ that for all $X,X^{\prime}$ differing only in $X_{G_{m}}$ (for any $m\in[M]$) that
$|\hat{V}_{k}(X)-\hat{V}_{k}(X^{\prime})|\leq C_{3}\frac{N^{2}}{M^{2}}(\log M+1)$
and it follows by McDiarmid’s inequality that
$P(|\hat{V}_{k}-\mathbb{E}\hat{V}_{k}|\geq\epsilon)\leq 2\exp\left(\frac{-2\epsilon^{2}}{M(CC_{1})^{2}N^{4}M^{-4}(\log M+1)^{2}}\right)$
which implies that $|V-\mathbb{E}\hat{V}_{k}|=O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right)$. ∎
@@
Recall that $C_{ij}$ is given by (128). To show that $\hat{V}_{k}$ is not anti-conservative, we observe that
$\mathbb{E}\hat{V}_{k}$ $=\sum_{i,j\in{\mathcal{I}}}\ \mathbb{E}\left[\mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg|T_{i},T_{j}\right]\right]$
$=\sum_{i,j\in{\mathcal{I}}}\ \mathbb{E}\left[C_{ij}(T_{i},T_{j})\cdot\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}\ \bigg|T_{i},T_{j}\right]\right]$
$=\sum_{i,j\in{\mathcal{I}}}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})\cdot C_{ij}(t_{i},t_{j})\cdot$
@@
$\geq\sum_{i,j\in{\mathcal{I}}}\ \sum_{t_{i},t_{j}}P(T_{i}=t_{i},T_{j}=t_{j})C_{ij}(t_{i},t_{j})$ (142)
$\geq\sum_{i,j\in{\mathcal{I}}}\ {\operatorname{Cov}}(v_{j},v_{j})$ (143)
where (142) holds by Lemma 3, and (143) holds by Lemma 4. It can be seen that (143) equals the claim of (67) made by Theorem 7. ∎
Proof of Theorem 7, unbiasedness claim.
To verify the conditions for unbiasedness given by Theorem 7, we observe that if $T_{i}$ is the constant mapping for all units, then $\hat{V}_{k}$ simplifies to
$\hat{V}_{k}=\sum_{i,j\in{\mathcal{I}}}{{\operatorname{Cov}}}\left(v_{i},v_{j}\right)\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)}$
where $Y_{ki}$ is constant for all $i\in[N]$ such that $i\in{\mathcal{I}}$. As this is an importance sampled estimator of covariance terms appearing in ${\operatorname{Var}\,}\tau_{k}^{{\textup{Haj}}}$, it is unbiased provided that the sampling probabilities $P\left(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right)$ (which equals $P(\bar{Z}_{i}=Y_{ki})$ for $i=j$) are non-zero for each $i,j\in{\mathcal{I}}$.
Proof of Lemmas 2 - 6
@@
For $i,j$ satisfying $i,j\in{\mathcal{I}}$, let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that
@@
Recall $C_{ij}$ is given by (128). To show (129), we observe that for $i,j\in\mathcal{J}_{0}$ that
@@
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]+\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]$
$={\operatorname{Cov}}\left(\frac{N}{\hat{N}_{Y_{ki}}},\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right)-\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
@@
$\hskip 28.45274pt-\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\right)\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]$
$\hskip 28.45274pt-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i,j\}}\right]\left(\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]\right)$
$\leq O\left(\frac{\log M}{M}\right)$ (145)
@@
To show (130), we observe for $i,j\in\mathcal{J}_{1}$ that
@@
$\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|F_{\{i,j\}}\right]-\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho}$ (146)
$\leq\frac{1}{\rho^{2}}\left(\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{ki}})^{2}}|F_{\{i,j\}}\right]\cdot\mathbb{E}\left[\frac{N^{2}}{(\hat{N}_{Y_{kj}})^{2}}|F_{\{i,j\}}\right]\right)^{1/2}+$
$\hskip 28.45274pt\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|F_{\{i\}}\right]\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{kj}}}|F_{\{j\}}\right]\cdot\frac{1}{\rho}$
$=O\left(1\right)+O\left(\frac{\log M}{M}\right)$ (147)
where (146) uses Lemma 5 and (55), and (147) uses Lemma 6 and the identity
@@
$=O\left(\frac{1}{M}\right)+O(1)$
which holds by Lemma 6 as well. ∎
@@
Recall that $v_{i}$ and $C_{ij}$ are respectively given by (34) and (128). If $i,j\in{\mathcal{I}}$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then
@@
To show (132), we observe that if $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$ then
@@
To show (133), we observe that if $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
@@
$\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right]$
where the final inequality uses (54). ∎
@@
$\hat{N}_{Q}$ $=\sum_{i\in Q}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ $\hat{N}_{\bar{S}}$ $=\sum_{i\in\bar{S}}\frac{1\{\bar{Z}_{j}=Y_{ki}\}}{P\left(\bar{Z}_{j}=Y_{ki}|T_{j}\right)}$ (154)
@@
From (138), it can be seen that changing $X_{G_{m}}$ for any $m\in[M]$ changes at most $\frac{dCN}{M}$ entries in $Q$, and hence changes $\hat{N}_{Q}$ by at most $\frac{dCN}{M\rho}$, where we have used (54). As a result, by McDiarmid’s inequality it holds for $C_{1}=dC/\rho$ that
$P\left(\left|\hat{N}_{Q}-N_{Q}\right|\geq\epsilon\ \Big|\ X_{\Gamma_{S}}\right)\leq\exp\left(-\frac{\epsilon^{2}M}{C_{1}^{2}N^{2}}\right)$ (155)
Thus for $\epsilon_{B}=\frac{C_{1}N}{\sqrt{M}}\sqrt{(2\delta^{-1}+1)\log M}$ and the indicator
@@
it follows from (155) that
$P(B=1\ |\ X_{\Gamma_{S}})$ $\leq\frac{1}{M^{\frac{2}{\delta}+1}}$
$\leq\frac{1}{N^{2+\delta}}$ (156)
where the last inequality uses that $|G_{m}|\leq N^{1-\delta}$, which implies $M\geq N^{\delta}$.
For $i,j\in{\mathcal{I}}$, we let $F_{\{i\}}$ and $F_{\{i,j\}}$ denote sigma-fields such that